# Paired kernels and truncated Toeplitz operators

### v1

@@ Abstract
Keywords: paired operator; paired kernel; Toeplitz operator; nearly-invariant subspace; truncated Toeplitz operator; kernel.
MSC (2010): 30H10, 47B35, 47B38
This work was partially supported by FCT/Portugal through CAMGSD, IST-ID, projects UIDB/04459/2020 and UIDP/04459/2020.
1 Introduction
@@
$S_{a,b}=aP^{+}+bP^{-}\quad\hbox{and}\quad\Sigma_{a,b}=P^{+}aI+P^{-}bI.$ (1.2)
In each case the pair $(a,b)$ is called the symbol pair. These operators have been studied by several mathematicians [5, 11, 13, 16, 17, 18, 22, 24, 26], often in the context of singular integral operators. If the operators $A$ and $B$ are invertible in $\mathcal{L}(X)$, then the paired operators are said to be of normal type [22], and this is the most studied case by far. Recently, several new results were obtained for operators of the form (1.2) with $a,b\in L^{\infty}$ satisfying weaker conditions [5, 11]:
$a\neq 0\hbox{ a.e.\ on }\mathbb{T},\quad b\neq 0\hbox{ a.e.\ on }\mathbb{T}$ (1.3)
@@
$\hbox{either}\quad a=b\quad\hbox{or}\quad a-b\neq 0\hbox{ a.e.\ on }\mathbb{T},$ (1.4)
which we assume in this paper. If $a\neq b$ we say that $(a,b)$ is nondegenerate.
Here we focus on paired operators of the form $S_{a,b}$ and their kernels. Paired operators are closely connected with Toeplitz operators. Indeed, on the one hand they are dilations of Toeplitz operators; on the other hand, if $b\in\mathcal{G}L^{\infty}$, then $S_{a,b}$ is equivalent after extension to the Toeplitz operator $T_{a/b}$ [3], implying that those two types of operator share many properties such as invertibility and Fredholmness, and their kernels are isomorphic. (See, for instance, [1] for the definition and properties of operators that are equivalent after extension.)
@@
$\mathop{\rm ker}\nolimits S_{a,b}\cong\mathop{\rm ker}\nolimits T_{a/b}$ (1.5)
@@
$P^{+}\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}.$ (1.6)
@@
$\mathop{\rm ker}\nolimits T_{\theta g}\subseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }\theta\hbox{ is an inner function},$ (1.7)
@@
$h_{+}\mathop{\rm ker}\nolimits T_{h_{+}g}\subseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }h_{+}\in H^{\infty},$ (1.8)
@@
$\mathop{\rm ker}\nolimits T_{h_{-}g}\supseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }h_{-}\in\overline{H^{\infty}},$ (1.9)
@@
$\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits S_{ah,bh}\quad\hbox{ for all }h\in L^{\infty}\setminus\{0\},$ (1.10)
@@
It can be shown, however, that the projection $P^{+}$ maps $\mathop{\rm ker}\nolimits S_{a,b}$ bijectively onto $P^{+}\mathop{\rm ker}\nolimits S_{a,b}={\rm ker}^{+}S_{a,b}$. Thus we study here the relations between the projected paired kernels with symbol pairs related by multiplication operators, for which one can obtain inclusions and equalities analogous to (1.7)–(1.9). These relations have also been sudied in [11], but we present them here in Section 3 for completeness, on the one hand, and also because several results are generalized. In Section 4 we apply them to study and describe the kernels of a class of asymmetric truncated Toeplitz operators (ATTO) of the form
$A^{\theta,\alpha}_{\phi}:K_{\theta}\to K_{\alpha},\qquad A^{\theta,\alpha}_{\phi}=P_{\alpha}\phi P_{\theta}{}_{|K_{\theta}},$ (1.11)
@@
$\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2_{+}}},\qquad A_{1-},B_{1-}\in\overline{H^{\infty}},\quad A_{2+},B_{2+}\in H^{\infty},$ (1.12)
@@
2 Projected paired kernels and near invariance
In this section we summarize some of the basic properties of projected paired kernels. A more detailed treatment is given in [11].
@@ 2 Projected paired kernels and near invariance
${\rm ker}^{-}S_{a,b}=\overline{z}\overline{{\rm ker}^{+}S_{\overline{b},\overline{a}}}.$ (2.1)
@@
$T:{\rm ker}^{+}S_{\overline{b},\overline{a}}\to{\rm ker}^{-}S_{a,b},\qquad\psi_{+}\mapsto\overline{z}\overline{\psi_{+}}$ (2.2)
@@
$M_{-a/b}:{\rm ker}^{+}S_{a,b}\to{\rm ker}^{-}S_{a,b},\qquad\phi_{+}\mapsto-\frac{a}{b}\phi_{+}.$ (2.3)
In general we omit the index $a/b$ and write simply $M=M_{-a/b}$. So for $\phi_{+}\in H^{2}_{+}$ we have
$\phi_{-}=M\phi_{+}\iff\phi_{-}\in H^{2}_{-}\quad\hbox{and}\quad a\phi_{+}+b\phi_{-}=0.$ (2.4)
@@
$f\in\mathcal{M},\ f(0)=0\implies S^{*}f\in\mathcal{M}.$ (2.5)
@@
${\rm ker}^{+}S_{a,b}\cap\theta H^{2}_{+}=\theta\,{\rm ker}^{+}S_{a\theta,b}.$ (2.6)
@@
$\phi_{+}\in{\rm ker}^{+}S_{a,b},\quad\overline{\theta}\phi_{+}\in H^{2}_{+}\iff\phi_{+}\in\theta\,{\rm ker}^{+}S_{a\theta,b},$ (2.7)
@@
${\rm ker}^{-}S_{a,b}\cap\overline{\theta}H^{2}_{-}=\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b},$ (2.8)
@@
$\phi_{-}\in{\rm ker}^{-}S_{a,b},\quad\theta\phi_{-}\in H^{2}_{-}\iff\phi_{-}\in\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b}.$ (2.9)
@@
$\phi_{+}\in\theta\,{\rm ker}^{+}S_{a\theta,b}\iff\phi_{-}\in{\rm ker}^{-}S_{a,\overline{\theta}b}$ (2.11)
@@
$\phi_{+}\in{\rm ker}^{+}S_{a\theta,b}\iff\phi_{-}\in\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b}.$ (2.12)
3 Inclusion relations
Let us now study the relations between two paired kernels with symbol pairs that are related by multiplication operators, by means of certain inclusion relations between their projections into $H^{2}_{\pm}$. Although some of these relations were obtained in [11] we present here all proofs for completeness, which also allows us in some cases to present alternative proofs.
@@ 3 Inclusion relations
${\rm ker}^{+}S_{a\theta_{1},b\overline{\theta}_{2}}\subseteq{\rm ker}^{+}S_{a,b},$ (3.1)
@@
${\rm ker}^{-}S_{a\theta_{1},b\overline{\theta}_{2}}\subseteq{\rm ker}^{-}S_{a,b},$ (3.2)
@@
${\rm ker}^{+}S_{a,bh_{-}}={\rm ker}^{+}S_{a,b}\quad\hbox{if}\quad\frac{a}{bh_{-}}\in L^{\infty},$ (3.3)
@@
${\rm ker}^{+}S_{ah_{-},b}={\rm ker}^{+}S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}\quad\hbox{if}\quad a/b\in L^{\infty};$ (3.4)
@@
${\rm ker}^{+}S_{gh_{-},h_{+}}=\mathop{\rm ker}\nolimits T_{gh_{+}^{-1}}=h_{+}\mathop{\rm ker}\nolimits T_{g}.$ (3.5)
@@
${\rm ker}^{+}S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b},\quad\hbox{if}\quad a/b\in L^{\infty},$ (3.6)
@@
(ii) if $h_{+}\in\mathcal{G}H^{\infty}$ or if $h_{+}$ is outer, $\dfrac{b}{ah_{+}}\in L^{\infty}$, then
$h_{+}{\rm ker}^{+}S_{ah_{+},b}={\rm ker}^{+}S_{a,b}\quad\hbox{and}\quad h_{+}{\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{a,bh_{+}};$
(iii) if $h_{+}$ is not outer, then the inclusions in (i) are strict.
@@
(ii) It is left to show that ${\rm ker}^{+}S_{a,b}\subseteq h_{+}{\rm ker}^{+}S_{ah_{+},b}$ and ${\rm ker}^{+}S_{a,bh_{+}}\subseteq h_{+}{\rm ker}^{+}S_{a,b}$. For the first inclusion, we have that
@@
If $h_{+}\in\mathcal{G}H^{\infty}$, then ${\phi_{+}}/{h_{+}}\in H^{2}_{+}$; if $\dfrac{b}{ah_{+}}\in L^{\infty}$, then $\phi_{+}/h_{+}\in L^{2}\cap\mathcal{N}_{+}$ and so $\phi_{+}/h_{+}\in H^{2}_{+}$. It follows that $\phi_{+}/h_{+}\in{\rm ker}^{+}S_{ah_{+},b}$.
On the other hand, for the second inclusion, we have $a\phi_{+}+(bh_{+})\phi_{-}=0\iff a\phi_{+}/h_{+}+b\phi=0$ and if $b/a\in L^{\infty}$ then $\phi_{+}/h_{+}\in L^{2}\cap\mathcal{N}_{+}=H^{2}_{+}$, so $\phi_{+}\in h_{+}\mathop{\rm ker}\nolimits S_{a,b}$.
(iii) If $h_{+}$ is not outer, we cannot have ${\rm ker}^{+}S_{a,b}\subseteq h_{+}{\rm ker}^{+}S_{ah_{+},b}\subseteq h_{+}H^{2}_{+}$, nor ${\rm ker}^{+}S_{a,bh_{+}}\subseteq h_{+}{\rm ker}^{+}S_{a,b}\subseteq h_{+}H^{2}_{+}$, by near invariance (Cor. 4.2 in [11]). ∎
@@
$h_{+}\mathop{\rm ker}\nolimits T_{a}\subseteq{\rm ker}^{+}S_{a,h_{+}}$ (3.7)
@@
$h_{-}{\rm ker}^{-}S_{a,bh_{-}}\subseteq{\rm ker}^{-}S_{a,b}\quad\hbox{and}\quad h_{-}{\rm ker}^{-}S_{a,b}\subseteq{\rm ker}^{-}S_{ah_{-},b}.$ (3.9)
@@
$\theta(z)=\exp\left(\frac{z-1}{z+1}\right),\qquad(z\in\mathbb{T}).$ (3.10)
@@
$(z+1)K_{\theta}=(z+1){\rm ker}^{+}S_{\overline{\theta},1}={\rm ker}^{+}S_{\overline{\theta},z+1}.$ (3.11)
@@
$(z+1)K_{\theta}={\rm ker}^{+}S_{\overline{\theta}\overline{z},(z+1)/z}\subseteq{\rm ker}^{+}S_{\overline{\theta}\overline{z},1}=\mathop{\rm ker}\nolimits T_{\overline{\theta}\overline{z}}=K_{\theta z},$ (3.12)
@@
${\rm ker}^{+}S_{\alpha B_{-},B_{+}}\quad\hbox{with}\quad B_{+},\overline{B_{-}}\in H^{\infty}.$ (3.13)
@@
${\rm ker}^{+}S_{a\theta,b}\subsetneq{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq{\rm ker}^{+}S_{a\theta_{2},b}\subsetneq\ldots\subsetneq{\rm ker}^{+}S_{a,b},$ (3.14)
@@
$\theta{\rm ker}^{+}S_{a\theta,b}\subsetneq\theta_{1}{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq\theta_{2}{\rm ker}^{+}S_{a\theta_{2},b}\subsetneq\ldots\subsetneq{\rm ker}^{+}S_{a,b},$ (3.15)
@@
$\overline{\theta}{\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq\theta_{1}{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq\theta_{2}{\rm ker}^{-}S_{a,\overline{\theta_{2}}b}\subsetneq\ldots\subseteq{\rm ker}^{-}S_{a,b},$ (3.16)
@@
${\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{2}}b}\subsetneq\ldots\subsetneq{\rm ker}^{-}S_{a,b}.$ (3.17)
@@
${\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{aB,b}+\mathop{\rm span}\nolimits\{\phi_{1+},\ldots,\phi_{k+}\}.$ (3.18)
@@
${\rm ker}^{+}S_{a,b}=B{\rm ker}^{+}S_{aB,b}+\mathop{\rm span}\nolimits\{\psi_{1+},\ldots,\psi_{k+}\}.$ (3.19)
@@
${\rm ker}^{+}S_{\alpha B_{-},B_{+}}={\rm ker}^{+}S_{\alpha\overline{\theta_{1}}\overline{\theta_{2}}\overline{O_{1}},O_{2}}=O_{2}\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}$ (3.20)

### v2

@@ Abstract.
Key words and phrases:
Paired operator, paired kernel, Toeplitz operator, nearly-invariant subspace, truncated Toeplitz operator, kernel
1991 Mathematics Subject Classification:
Primary 30H10; Secondary 47B35, 47B38
@@
1. Introduction
@@
(1.2) $S_{a,b}=aP^{+}+bP^{-}\quad\hbox{and}\quad\Sigma_{a,b}=P^{+}aI+P^{-}bI.$
In each case the pair $(a,b)$ is called the symbol pair. These operators have been studied by several mathematicians [5, 11, 13, 16, 17, 18, 22, 24, 26], often in the context of singular integral operators. If the operators $A$ and $B$ are invertible in $\mathcal{L}(X)$, then the paired operators are said to be of normal type [22], and this is the most studied case by far.
Recently, several new results were obtained for operators of the form (1.2) with $a,b\in L^{\infty}$ satisfying weaker conditions [5, 11]:
(1.3) $a\neq 0\hbox{ a.e.\ on }\mathbb{T},\quad b\neq 0\hbox{ a.e.\ on }\mathbb{T}$
@@
(1.4) $\hbox{either}\quad a=b\quad\hbox{or}\quad a-b\neq 0\hbox{ a.e.\ on }\mathbb{T},$
which will be standing assumptions in this paper. If $a\neq b$ we say that $(a,b)$ is nondegenerate.
Here we focus on paired operators of the form $S_{a,b}$ and their kernels. Paired operators are closely connected with Toeplitz operators. Indeed, on the one hand they are dilations of Toeplitz operators; on the other hand, if $b\in\mathcal{G}L^{\infty}$ (that is, if $b$ is invertible in $L^{\infty}$), then $S_{a,b}$ is equivalent after extension to the Toeplitz operator $T_{a/b}$ [3], implying that those two types of operator share many properties such as invertibility and Fredholmness, and their kernels are isomorphic. (See, for instance, [1] for the definition and properties of operators that are equivalent after extension.)
@@
(1.5) $\mathop{\rm ker}\nolimits S_{a,b}\cong\mathop{\rm ker}\nolimits T_{a/b}$
@@
(1.6) $P^{+}\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}.$
@@
(1.7) $\mathop{\rm ker}\nolimits T_{\theta g}\subseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }\theta\hbox{ is an inner function},$
@@
(1.8) $h_{+}\mathop{\rm ker}\nolimits T_{h_{+}g}\subseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }h_{+}\in H^{\infty},$
@@
(1.9) $\mathop{\rm ker}\nolimits T_{h_{-}g}\supseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }h_{-}\in\overline{H^{\infty}},$
@@
[…]
@@
It can be shown, however, that under our standing assumptions (1.3) and (1.4) the projection $P^{+}$ maps $\mathop{\rm ker}\nolimits S_{a,b}$ bijectively onto $P^{+}\mathop{\rm ker}\nolimits S_{a,b}={\rm ker}^{+}S_{a,b}$, indeed $P^{+}\phi=\dfrac{b}{b-a}\phi$, given that (1.4) holds (see [11, Prop. 2.1] for further details).
Thus we study here the relations between the projected paired kernels with symbol pairs related by multiplication operators, for which one can obtain inclusions and equalities analogous to (1.7)–(1.9). These relations have also been sudied in [11], but we present them here in Section 3 for completeness, on the one hand, and also because several results are generalized. In Section 4 we apply them to study and describe the kernels of a class of asymmetric truncated Toeplitz operators (ATTO) of the form
(1.11) $A^{\theta,\alpha}_{\phi}:K_{\theta}\to K_{\alpha},\qquad A^{\theta,\alpha}_{\phi}=P_{\alpha}\phi P_{\theta}{}_{|K_{\theta}},$
@@
(1.12) $\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2_{+}}},\qquad A_{1-},B_{1-}\in\overline{H^{\infty}},\quad A_{2+},B_{2+}\in H^{\infty},$
@@
2. Projected paired kernels and near invariance
In this section we summarize some of the basic properties of projected paired kernels. A more detailed treatment is given in [11]. We remind the reader of our standing assumptions (1.3) and (1.4).
@@ 2. Projected paired kernels and near invariance
(2.1) ${\rm ker}^{-}S_{a,b}=\overline{z}\overline{{\rm ker}^{+}S_{\overline{b},\overline{a}}}.$
@@
(2.2) $T:{\rm ker}^{+}S_{\overline{b},\overline{a}}\to{\rm ker}^{-}S_{a,b},\qquad\psi_{+}\mapsto\overline{z}\overline{\psi_{+}}$
@@
(2.3) $M_{-a/b}:{\rm ker}^{+}S_{a,b}\to{\rm ker}^{-}S_{a,b},\qquad\phi_{+}\mapsto-\frac{a}{b}\phi_{+}$
(this is well-defined under the standing assumption (1.3)). In general we omit the index $a/b$ and write simply $M=M_{-a/b}$. So for $\phi_{+}\in H^{2}_{+}$ we have
(2.4) $\phi_{-}=M\phi_{+}\iff\phi_{-}\in H^{2}_{-}\quad\hbox{and}\quad a\phi_{+}+b\phi_{-}=0.$
@@
(2.5) $f\in\mathcal{M},\ f(0)=0\implies S^{*}f\in\mathcal{M}.$
@@
(2.6) ${\rm ker}^{+}S_{a,b}\cap\theta H^{2}_{+}=\theta\,{\rm ker}^{+}S_{a\theta,b}.$
@@
(2.7) $\phi_{+}\in{\rm ker}^{+}S_{a,b},\quad\overline{\theta}\phi_{+}\in H^{2}_{+}\iff\phi_{+}\in\theta\,{\rm ker}^{+}S_{a\theta,b},$
@@
(2.8) ${\rm ker}^{-}S_{a,b}\cap\overline{\theta}H^{2}_{-}=\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b},$
@@
(2.9) $\phi_{-}\in{\rm ker}^{-}S_{a,b},\quad\theta\phi_{-}\in H^{2}_{-}\iff\phi_{-}\in\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b}.$
@@
(2.11) $\phi_{+}\in\theta\,{\rm ker}^{+}S_{a\theta,b}\iff\phi_{-}\in{\rm ker}^{-}S_{a,\overline{\theta}b}$
@@
(2.12) $\phi_{+}\in{\rm ker}^{+}S_{a\theta,b}\iff\phi_{-}\in\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b}.$
3. Inclusion relations
Let us now study the relations between two paired kernels with symbol pairs that are related by multiplication operators, by means of certain inclusion relations between their projections into $H^{2}_{\pm}$. Although some of these relations were obtained in [11] we present here all proofs for completeness, which also allows us in some cases to present alternative proofs. We again remind the reader of our standing assumptions (1.3) and (1.4).
@@ 3. Inclusion relations
(3.1) ${\rm ker}^{+}S_{a\theta_{1},b\overline{\theta}_{2}}\subseteq{\rm ker}^{+}S_{a,b},$
@@
(3.2) ${\rm ker}^{-}S_{a\theta_{1},b\overline{\theta}_{2}}\subseteq{\rm ker}^{-}S_{a,b},$
@@
(3.3) ${\rm ker}^{+}S_{a,bh_{-}}={\rm ker}^{+}S_{a,b}\quad\hbox{if}\quad\frac{a}{bh_{-}}\in L^{\infty},$
@@
(3.4) ${\rm ker}^{+}S_{ah_{-},b}={\rm ker}^{+}S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}\quad\hbox{if}\quad a/b\in L^{\infty};$
@@
(3.5) ${\rm ker}^{+}S_{gh_{-},h_{+}}=\mathop{\rm ker}\nolimits T_{gh_{+}^{-1}}=h_{+}\mathop{\rm ker}\nolimits T_{g}.$
@@
(3.6) ${\rm ker}^{+}S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b},\quad\hbox{if}\quad a/b\in L^{\infty},$
@@
(ii) if $h_{+}\in\mathcal{G}H^{\infty}$, or if $h_{+}$ is outer with $\dfrac{b}{ah_{+}}\in L^{\infty}$, then
$h_{+}{\rm ker}^{+}S_{ah_{+},b}={\rm ker}^{+}S_{a,b};$
(iii) if $h_{+}\in\mathcal{G}H^{\infty}$, or if $h_{+}$ is outer with $\dfrac{b}{a}\in L^{\infty}$, then
$h_{+}{\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{a,bh_{+}};$
(iv) if $h_{+}$ is not outer, then the inclusions in (i) are strict.
@@
(ii) It is left to show that ${\rm ker}^{+}S_{a,b}\subseteq h_{+}{\rm ker}^{+}S_{ah_{+},b}$. We have that
@@
If $h_{+}\in\mathcal{G}H^{\infty}$, then ${\phi_{+}}/{h_{+}}\in H^{2}_{+}$; if $h_{+}$ is outer and $\dfrac{b}{ah_{+}}\in L^{\infty}$, then $\phi_{+}/h_{+}\in L^{2}\cap\mathcal{N}_{+}$ and so $\phi_{+}/h_{+}\in H^{2}_{+}$. It follows that $\phi_{+}/h_{+}\in{\rm ker}^{+}S_{ah_{+},b}$.
(iii) Now it remains to show that ${\rm ker}^{+}S_{a,bh_{+}}\subseteq h_{+}{\rm ker}^{+}S_{a,b}$. We have
$a\phi_{+}+(bh_{+})\phi_{-}=0\iff a\phi_{+}/h_{+}+b\phi_{-}=0.$
If $h_{+}\in\mathcal{G}H^{\infty}$, then $\phi_{+}/h_{+}\in H^{2}_{+}$; if $h_{+}$ is outer and $b/a\in L^{\infty}$ then $\phi_{+}/h_{+}\in L^{2}\cap\mathcal{N}_{+}=H^{2}_{+}$, so $\phi_{+}\in h_{+}{\rm ker}^{+}S_{a,b}$.
(iv) If $h_{+}$ is not outer, we cannot have ${\rm ker}^{+}S_{a,b}\subseteq h_{+}{\rm ker}^{+}S_{ah_{+},b}\subseteq h_{+}H^{2}_{+}$, nor ${\rm ker}^{+}S_{a,bh_{+}}\subseteq h_{+}{\rm ker}^{+}S_{a,b}\subseteq h_{+}H^{2}_{+}$, by near invariance (Cor. 4.2 in [11]). ∎
@@
(3.7) $h_{+}\mathop{\rm ker}\nolimits T_{a}\subseteq{\rm ker}^{+}S_{a,h_{+}}$
@@
(3.9) $h_{-}{\rm ker}^{-}S_{a,bh_{-}}\subseteq{\rm ker}^{-}S_{a,b}\quad\hbox{and}\quad h_{-}{\rm ker}^{-}S_{a,b}\subseteq{\rm ker}^{-}S_{ah_{-},b}.$
@@
(3.10) $\theta(z)=\exp\left(\frac{z-1}{z+1}\right),\qquad(z\in\mathbb{T}).$
@@
(3.11) $(z+1)K_{\theta}=(z+1){\rm ker}^{+}S_{\overline{\theta},1}={\rm ker}^{+}S_{\overline{\theta},z+1}.$
@@
(3.12) $(z+1)K_{\theta}={\rm ker}^{+}S_{\overline{\theta}\overline{z},(z+1)/z}\subseteq{\rm ker}^{+}S_{\overline{\theta}\overline{z},1}=\mathop{\rm ker}\nolimits T_{\overline{\theta}\overline{z}}=K_{\theta z},$
@@
(3.13) ${\rm ker}^{+}S_{\alpha B_{-},B_{+}}\quad\hbox{with}\quad B_{+},\overline{B_{-}}\in H^{\infty}.$
@@
(3.14) ${\rm ker}^{+}S_{a\theta,b}\subsetneq{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq{\rm ker}^{+}S_{a\theta_{2},b}\subsetneq\ldots\subsetneq{\rm ker}^{+}S_{a,b},$
@@
(3.15) $\theta{\rm ker}^{+}S_{a\theta,b}\subsetneq\theta_{1}{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq\theta_{2}{\rm ker}^{+}S_{a\theta_{2},b}\subsetneq\ldots\subsetneq{\rm ker}^{+}S_{a,b},$
@@
(3.16) $\overline{\theta}{\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq\theta_{1}{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq\theta_{2}{\rm ker}^{-}S_{a,\overline{\theta_{2}}b}\subsetneq\ldots\subseteq{\rm ker}^{-}S_{a,b},$
@@
(3.17) ${\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{2}}b}\subsetneq\ldots\subsetneq{\rm ker}^{-}S_{a,b}.$
@@
(3.18) ${\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{aB,b}+\mathop{\rm span}\nolimits\{\phi_{1+},\ldots,\phi_{k+}\}.$
@@
(3.19) ${\rm ker}^{+}S_{a,b}=B{\rm ker}^{+}S_{aB,b}+\mathop{\rm span}\nolimits\{\psi_{1+},\ldots,\psi_{k+}\}.$
@@
(3.20) ${\rm ker}^{+}S_{\alpha B_{-},B_{+}}={\rm ker}^{+}S_{\alpha\overline{\theta_{1}}\overline{\theta_{2}}\overline{O_{1}},O_{2}}=O_{2}\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}$